OpenStudy (anonymous):

A man uses a lever jack to lift a 2.00 * 10^3 kg car to change a tire. Each stroke of the jack lift the car a distance of 5.00 * 10^-3 m. The vertical displacement of the handle is 0.75 m for each stroke. How much total force does he use with each stroke of the jack? A. 3.0 * 10^5 N B. 2.9 * 10^6 N C. 131 N D. 19, 600 N

OpenStudy (michele_laino):

here we have to apply the virtual work principle, namely the motor work has to be equal to the resistive work. So we can write: \[F \cdot 5 \times {10^{ - 3}} = 2 \times {10^3} \times 0.75\] what is F?

OpenStudy (michele_laino):

oops.. I have made a typo, here is the right formula: \[\Large F \cdot 5 \times {10^{ - 3}} = 2 \times {10^3} \times 9.81 \times 0.75\]

OpenStudy (michele_laino):

since we know the mass of our car, which is 2,000 Kg

OpenStudy (anonymous):

2943000 ? so our solution is choice B?

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

yay!! thank you!! :D

OpenStudy (michele_laino):

:)